OpenStudy (flamingpiano):

Can someone help me with this? 2 log2 2 + 2 log2 6 − log2 3x = 3 is the equation. I'm new to logs, and I just tried out the problem. Here's what I did... 2 log2 2 + 2 log2 6 − log2 3x = 3 log2 4 + log2 36 - log2 3x = 3 4 + 36 - 3x = 3 40 - 3x = 3 -3x = -37 x = 12 1/3 However... I know this is incorrect. Since the answer choices are x = 2 x = 6 x = 16 x = 18 Please don't give me the answer, I need to know how to do it.

OpenStudy (anonymous):

\begin{bmatrix}\mathrm{Solution:}\:&\:-\infty \:<f\left(x\right)<\infty \\\:\mathrm{Interval\:Notation:}&\:\left(-\infty \:,\:\infty \:\right)\end{bmatrix}

OpenStudy (flamingpiano):

Uh, I didn't see that in the lesson. What do I do with the ∞?

OpenStudy (anonymous):

whoops wrong 1 sorry

OpenStudy (anonymous):

can u type the equation correct plz

OpenStudy (anonymous):

144 / 3x = 2^3 = 8 144/8 = 3x 18 = 3x

OpenStudy (anonymous):

here this should help bro

OpenStudy (anonymous):

step 1) use the power formula of logarithms: x logb (a) = logb a^x: so, 2 log2 (2) + 2 log2 (6) - log2 (3x) = 3 becomes log2(2^2) + log2 (6^2) - log2 (3x) = 3 step 2) simplify: log2 (4) + log2 (36) - log2 (3x) = 3 step 3) use the product rule of logarithms: logb (a) + logb (c) = logb (ac). So, log2 (4) + log2 (36) - log2 (3x) = 3 becomes log2 (36*4) - log2 (3x) =3 step 4) simplify: log2 (144) - log2 (3x) = 3 step 5) use the quotient rule of logarithms: logb (a) - logb (c) = logb (a/c). So, log2 (144) - log2 (3x) = 3 becomes log2 (144/3x) = 3 step 6) transform the logarithmic equation into an exponential one: logb (x) = y is the same as b^y=x So, log2 (144/3x) = 3 becomes 2^3 = 144/3x step 7) Simplify: 8 = 144/3x step 8) multiply both sides by 3x: 3x*8 = (144/3x)*3x step 9) simplify: 24x = 144 step 10) divide both sides by 24 and simplify: 24x/24 = 144/24 the answer should be 6.

OpenStudy (anonymous):

does this help u any bro

OpenStudy (flamingpiano):

@usercode3rror Yeah, thanks! Sorry, my internet went out for a bit.